Given the following enthalpy change values for the reactions below:

A + 2B → C ; ΔrH = -228.3 kJ mol-1

E + F → 2C + D ; ΔrH = 493.9 kJ mol-1

½ E → G + 2B ; ΔrH = 314.3 kJ mol-1

Calculate ΔrH for the following reaction (in kJ mol-1) using Hess's Law:

2A + D → F + 2G
1
Expert's answer
2019-10-21T07:15:55-0400

A + 2B =C.......1


E + F = 2C + D........2

1/2 E = G + 2B........3

According to hess law if we do (equation no 1 × 2) - ( equation no 2) + (equation no 3)

In the result we will reach at 2A + D = F + 2G

So ∆rH of these reaction will be -636.2 KJ mol-1